A weak acid HA has Ka=10^-6. What would be the molar ratio of this acid and its salt with strong base so that the pH of the buffer solution is 5?
1
Expert's answer
2016-03-09T08:36:43-0500
The formula of the pH of buffer solution:
pH = pKa(acid)+lg(C(salt)/C(acid))
5 = -Lg(10-6) + lg(C(salt)/C(acid))
5 = 6 + lg(C(salt)/C(acid))
-1 = lg(C(salt)/C(acid))
10-1 = C(salt)/C(acid)
0.1 = C(salt)/C(acid)
C(acid)/ C(salt) = 10

Answer:
The molar ratio of the acid and its salt with strong base is 10.